OpenStudy (anonymous):

If u = 8 + 4i and v = 3i + 6, then what is u – v ? a)2+2i b)5-2i C)1+i D)2+i E)2-i

OpenStudy (anonymous):

@leoh1998

OpenStudy (michele_laino):

by definition, we have: \[\Large \begin{gathered} u - v = \left( {8 + 4i} \right) - \left( {6 + 3i} \right) = 8 + 4i - 6 - 3i = \hfill \\ \hfill \\ = \left( {8 - 6} \right) + i\left( {4 - 3} \right) = ...? \hfill \\ \end{gathered} \]